# Forest-Chat: Adapting Vision-Language Agents for Interactive Forest Change Analysis

### v1

@@ 4.1 Multi-level Change Interpretation Model
The MCI model is a dual-branch VLM designed to process pairs of bi-temporal RS images, as shown in Figure 5(a). It employs a Siamese SegFormer B1 backbone (Xie et al., 2021) to extract multi-scale semantic features from each image, providing both low-level detail and high-level contextual information. These features form the shared representation from which two task-specific branches are derived.
Bi-temporal Iterative Interaction (BI3) layer: A central innovation of the MCI model that is embedded within the change detection and change captioning branches, which aims to enhance and fuse bi-temporal features. It combines two modules: Local Perception Enhancement (LPE), and Global Difference Fusion Attention (GFDA) to extract discriminative features of interest.
The LPE module applies convolutions of different kernel shapes to capture variations across multiple spatial scales, producing feature maps that enrich the diversity of feature information and the ability to detect local changes. Subsequently, the GDFA module uses feature differencing and attention to highlight regions of true change while down-weighting irrelevant noise. Following the LPE and GDFA modules, Layer Normalisation (LN) and a Multi-Layer Perceptron (MLP) are applied to the bi-temporal features to enhance them. Stacking BI3 layers across the branches progressively strengthens feature discrimination. This implementation uses a total of three BI3 layers per branch.
Change detection branch (top right, Fig. 5(a)): Applies stacked BI3 layers with residual connections to iteratively enhance and refine the high-level bi-temporal features extracted from the backbone. This improves the model’s capacity to capture meaningful semantic differences between time points. To recover finer spatial detail, the branch then employs Convolution-Based Bi-temporal Fusion (CBF) modules across four spatial scales to give four window sizes, progressively integrating features from both images. Finally, a deconvolution pathway merges these multi-scale representations from bottom to top, enabling the production of precise, high-resolution change masks.
Change captioning branch (bottom right, Fig. 5(a)): Addresses semantic-level interpretation, transforming vision differences into natural language descriptions. It begins with multiple stacked BI3 layers, which refine and align high-level semantic features from the bi-temporal images. These enhanced features are then passed through a convolution-based projection layer, bridging the visual and textual domains. Finally, the projected features are decoded by a Transformer decoder operating with a 512-dimensional word embedding space, generating descriptive sentences of the observed changes.
@@ 4.2 LLM Agent Task Orchestration
The LLM functions as the controller of Change-Agent (Fig. 1 right side), directing task execution and dynamically responding to user queries. While some recent LLMs are equipped with native visual perception (openai et al., 2023), these capabilities are not yet optimally suited to RS imagery, which requires specialised feature extraction (Zhang and Wang, 2024). To address this, the system is extended with Python-based tools, including the backbone feature extractor, change detection and change captioning modules, and standard processing libraries. By generating and executing Python programs, the LLM can invoke these tools as needed, enabling automated multistep workflows without human intervention. Within the application interface, users are able to apply and remove tools as desired, including access to Google Scholar and Google Search for external knowledge. It is within the change perception module that the AnyChange (Zheng et al., 2024) model and additional forest change specific functions are made available to the central LLM, realising the idea of Forest-Chat.
In this framework, ChatGPT-4o-mini serves as the LLM backbone, selected to balance performance, speed, and cost. The system is capable of incorporating alternative LLMs (e.g. InternLM2.5-7B-Chat (Cai et al., 2024)) for task orchestration and execution, with selection predominantly based on the degree of complex reasoning required or availability of computing power. After processing tool outputs, the LLM synthesises results into natural language responses, whether as change masks, descriptive captions, object counts, deforestation percentages, predictive assessments, and causal reasoning. System prompts with few-shot examples are provided to the LLM to support forest change analysis workflows, ensuring reliable tool invocation and consistent output formatting. This enables the combination of detailed and grounded visual analysis with flexible reasoning and user-facing interaction.
@@ 4.3 AnyChange for Zero-shot Change Detection
The AnyChange model developed by Zheng et al. (2024) introduces a training-free framework for zero-shot change detection for RS imagery (Fig. 1 lower left, Fig. 5(b)). The method builds on the Segment Anything Model (SAM) (Kirillov et al., 2023), adapting its latent space for bi-temporal analysis without requiring fine-tuning. The model is designed to generalise across unseen change types and data distributions, producing either pixel-level or instance-level predictions. It supports three modes of operation: fully-automatic, semi-automatic (threshold-based), and prompt-guided using point inputs.
The approach is grounded in the structured properties of SAM’s latent space. Mask embeddings generated by the image encoder provide semantic information such that objects of the same category (e.g. water, buildings, bare land) cluster together within a single image. Furthermore, corresponding object categories across temporally separated images of the same location remain closely aligned despite differences in imaging conditions (e.g. seasonality). These intra-image and inter-image semantic consistencies provide the foundation for adapting SAM to RSCD without retraining or architectural modifications.
AnyChange employs a bi-temporal latent matching strategy in which object masks are first generated for each temporal image using SAM, and embeddings for each mask are computed by averaging the corresponding image encoder inputs. The semantic similarity between embeddings is quantified using a cosine-based metric, allowing changes to be identified as deviations between corresponding object embeddings over time. The method ensures temporal symmetry by performing bidirectional matching between time points. Change predictions can then be obtained either by ranking confidence scores or applying a threshold on embedding angles.
To enable object-centric change detection, AnyChange combines its bi-temporal latent matching with SAM’s point prompt mechanism, yielding a point query mechanism. Users provide a set of spatial coordinates on a single-temporal image, which are treated as point prompts corresponding to a specific category. SAM generates object proposals for these points, and the average mask embeddings are computed. These embeddings are then matched with proposals from both temporal images using cosine similarity, producing change scores that can be filtered using a user-defined angle threshold provided by the bi-temporal latent matching approach.
Instead of inputting a list of point coordinates, this paper develops an interactive application interface that allows users to click directly on areas of interest to generate point queries, improving convenience over manual coordinate entry.
@@ 4.4 Multi-task Loss Balancing Strategy
During the training of the MCI model, multi-task learning is performed on both change detection and change captioning. For both tasks, standard cross-entropy loss is used to quantify the dissimilarity between predicted change masks and ground truth annotations, and the dissimilarity between predicted sentences and their corresponding ground truth sentences respectively. The two losses have significant discrepancies between their magnitudes, which negatively impacts overall change interpretation performance, due to the risk of one task dominating the training signal of the other. A common approach, which is used by Liu et al. (2024), and consequently adopted here, is to normalise the losses to the same order of magnitude and detach their gradients for stability. This ensures that each task contributes equally to the final loss function, and enables a simultaneous optimisation of the dual change detection and captioning tasks. The final loss function is derived as:
@@ 4.5 Evaluation Metrics
To evaluate the performance of Forest-Chat’s supervised and zero-shot settings for binary change detection in remote sensing images, the Mean Intersection over Union (MIoU) metric is employed. MIoU quantifies segmentation quality by measuring the average overlap between predicted change masks and ground-truth masks across all classes. MIoU is defined below as:
@@
where $C$ is the number of classes (i.e. change (c), no change (nc) for which metrics are calculated. For each class i, $TP_{i}$ are true positives, $FP_{i}$ false positives, and $FN_{i}$ false negatives for class $i$.
By analysing both the individual class performances through IoU, as well as MIoU, model predictions for the under-represented change class can be compared. Overall accuracy (OA) is naturally a poor performance indicator for this task due to the severe class imbalances present in both datasets. Despite LEVIR-MCI-Trees containing three classes (i.e. road, building, background), performance metrics are calculated on the basis of change (c) and no change (nc), harmonising the results with those obtained for Forest-Change.
For evaluating the change captioning capabilities of FC-Supervised, the evaluation metrics employed by Liu et al. (2024) are re-used here for consistency. The four metrics are redefined here:
@@ 5.2 Forest-Chat Supervised Performance
There remains a lack of available methods that provide simultaneously capabilities for joint change detection and captioning (CDC). Consequently, performance evaluations are done by comparing FC-Supervised to SOTA models for the two tasks separately on the LEVIR-MCI-Trees and Forest-Change datasets. Change3D (Zhu et al., 2025) is capable of both tasks, but is not implemented in a unified manner. BiFA (Zhang et al., 2024a) is selected for change detection, and Chg2Cap (Chang and Ghamisi, 2023) for change captioning. Additionally, identifying differences in difficulty between the two datasets can be addressed, comparing performances in a low-data forest change analysis scenario against a more urban focused one constructed from a larger and more varied dataset.
Table 1 provides the performance comparisons for the benchmarked models on both datasets. For both datasets, FC-Supervised is the strongest performer for change captioning, and is the best on LEVIR-MCI-Trees for change detection. On the Forest-Change dataset, FC-Supervised is marginally outperformed by BiFA Zhang et al. (2024a) for change class identification, with both models achieving similar overall performance. These results demonstrate strong generalisation capabilities across varying data domains for both change detection and captioning. Among the benchmarked methods, FC-Supervised is the slowest at captioning inference and the second slowest for change detection (see Table 2). This is expected because it performs joint multi-task reasoning with heavy cross-modal and cross-temporal interactions and multi-task attention. However, these inference times are not prohibitive for practical applications, especially given that only a single model is required to generate pixel-level and semantic-level information of changes within RS image pairs. Zero-shot inference times are discussed in the next section.
@@
LEVIR-MCI-Trees BiFA 87.54 95.63 79.45 - - - - - - -
@@
Forest-Change BiFA 67.34 95.85 38.84 - - - - - - -
@@
Table 2: Inference time per iteration (s), averaged across both datasets.
@@
Change Captioning Change3D 4.00
@@
Change Detection Change3D 0.53
@@
FC-Zero-shot 31.88
@@
Overall change detection performance is considerably better for building and road detection than for deforestation, likely due to the pronounced class imbalances within Forest-Change, fuzzy edges at the borders of many deforestation patches, the uniqueness of deforestation patterns, and the relatively small dataset size. Within the Forest-Change dataset, a considerable percentage of change masks are comprised of numerous tiny patches that are challenging for models to detect. Correspondingly, this resulted in all benchmarked CD models frequently predicting change masks that mirrored the deforestation patterns observed, but contained few overlaps with the ground-truth masks (see row 4 in Fig. 6). Models generally perform well on larger patches, but would still struggle with accurately predicting fuzzy boundaries. This corroborates the findings of Liu et al. (2024), who highlight the segmentation performance of objects containing fewer than 400 pixels is low. Compared with deforestation detection, building change detection patterns and object geometries are more predictable, further raising the challenge posed by the Forest-Change dataset.
@@
Captioning performance evaluation is more nuanced. For the LEVIR-MCI-Trees dataset, the BLEU-1, BLEU-2, and CIDEr-D metrics are generally higher, while BLEU-3, BLEU-4, METEOR, and ROUGEL are better for Forest-Change. The observed results largely reflect the differences in annotation styles between the two datasets, as described in Section 3.2.3. Despite models frequently predicting deforestation severity accurately, they occasionally get confused about the location and characterisation of deforestation patches, and are unable to recognise geographic features within the imagery. Improved caption generation, or domain-aware approaches, such as vocabulary expansion (Gao et al., 2024) or domain-adaptive fine-tuning (Guo et al., 2022), could mitigate this by elevating the representation of geophysical descriptors in the latent space of these models.
5.3 Forest-Chat Zero-shot Performance
In keeping with the previous section, FC-Zero-shot performance is evaluated on both the Forest-Change and LEVIR-MCI-Trees datasets. To optimise performance, a Bayesian hyperparameter search comprising 20 runs was conducted to identify the most influential parameters. Following the Segment Any Change framework (Zheng et al., 2024), bi-temporal matching and feature normalisation are enabled throughout due to their consistently positive impact on change discrimination.

### v2

@@ 4.1 Multi-level Change Interpretation Model
The MCI model is a dual-branch VLM designed to process pairs of bi-temporal RS images, as shown in Figure 5(a). A Siamese SegFormer B1 backbone (Xie et al., 2021) extracts multi-scale semantic features from each image, providing both low-level detail and high-level contextual information, forming the shared representation from which two task-specific branches are derived.
The central innovation is the Bi-temporal Iterative Interaction (BI3) layer, embedded within both branches to enhance and fuse bi-temporal features through Local Perception Enhancement (LPE), and Global Difference Fusion Attention (GFDA) to extract discriminative features of interest. LPE applies multi-scale convolutions to capture local change variations, while GDFA uses feature differencing and attention to highlight regions of true change while suppressing noise. Three BI3 layers are stacked to progressively strengthen feature discrimination across both branches.
The Change detection branch (top right, Fig. 5(a)) applies stacked BI3 layers with residual connections, before employing Convolution-based Bi-temporal Fusion (CBF) modules across four spatial scales and a deconvolution pathway to produce precise, high-resolution change masks. The Change captioning branch (bottom right, Fig. 5(a)) refines semantic features through stacked BI3 layers, projects them into the textual domain via a convolution-based projection layer, and decodes them using a transformer decoder to generate natural language descriptions of observed changes.
@@ 4.2 LLM Agent Task Orchestration
The LLM functions as the controller of Forest-Chat (Fig. 1 right side), interpreting user instructions, selecting appropriate perception modules, directing task execution, and synthesising results into natural language responses. While some recent LLMs offer native visual perception (openai et al., 2023; Hurst et al., 2024), these capabilities are not yet optimally suited to RS imagery (Zhang and Wang, 2024), motivating the integration of specialised external tools including the MCI model, AnyChange, GPT-4o, and forest-specific analytical functions. By generating and executing Python programs, the LLM can invoke these tools as needed, enabling automated multistep workflows without human intervention. System prompts with few-shot examples, retry mechanisms, and output validation are provided to support forest change analysis workflows, ensuring reliable tool invocation and consistent output formatting.
ChatGPT-4o-mini serves as the LLM backbone, selected to balance performance, speed, and cost. GPT-4 models have demonstrated strong performance in agentic workflows over open-source models (Chang et al., 2024), with GPT-4o-mini retaining competitive performance at a fraction of the inference cost (Xiao et al., 2024). The framework is not restricted to any specific LLM, and alternative models (e.g. InternLM2.5-7B-Chat (Cai et al., 2024)) may be substituted based on reasoning ability, code generation capability, and available compute.
@@ 4.3 AnyChange for Zero-shot Change Detection
AnyChange (Zheng et al., 2024) introduces a training-free framework for zero-shot change detection in RS imagery (Fig. 1 lower left, Fig. 5(b)), building on the Segment Anything Model (SAM) (Kirillov et al., 2023) without requiring fine-tuning or architectural modifications. It exploits the structured semantic properties of SAM’s latent space to perform bi-temporal latent matching, comparing per-mask embeddings across temporal images using cosine similarity to identify semantically inconsistent regions. Temporal symmetry is ensured through bidirectional matching, with change predictions obtained by ranking confidence scores or thresholding on embedding angles. The model is designed to generalise across unseen change types and data distributions, producing either pixel-level or instance-level predictions. The model supports fully-automatic, semi-automatic (threshold-based), and point prompt-guided modes of operation.
To enable object-centric change detection, AnyChange combines bi-temporal latent matching with SAM’s point prompt mechanism. This yields a point query mechanism in which user-provided spatial coordinates are used to generate category-specific object proposals that are matched across time (Zheng et al., 2024). Rather than requiring manual coordinate entry, we develop an interactive interface to allow users to click directly on areas of interest to generate point queries. The effect of point prompts on improving performance for a selected semantic class is well documented in Zheng et al. (2024) and is not explored further here; for binary change detection, point prompts provide limited benefit given that change occurs within a single semantic class.
4.4 GPT-4o for Zero-shot Change Captioning and
Refinement
Forest-Chat incorporates GPT-4o (Hurst et al., 2024) for zero-shot change captioning and refinement, selected for its strong performance on single-image RS captioning (Ibrahim et al., ) and accessible multimodal API. Its application for RS change captioning remains largely unexplored (Soni et al., 2025), though open-source alternatives also warrant further investigation (León et al., 2026).
Bi-temporal image pairs are denormalised, base64-encoded, and passed to GPT-4o via the OpenAI multimodal API alongside a system prompt establishing an expert RS analyst role. Three prompting modes are evaluated. The general prompt instructs the model to describe changes concisely in a single sentence. The style-guided prompt provides few-shot caption examples and explicit style constraints to align outputs with the target caption distribution. The refinement prompt operates as a two-stage pipeline in which FC-Supervised first generates a predicted caption, which is then passed alongside the bi-temporal image pair to GPT-4o, tasked with preserving the core change vocabulary while enriching the description with spatial grounding and domain-specific context directly observable in the imagery. A temperature of 0.2 is used across all modes to reduce output variability. A summary of the prompting strategies is provided in A. Full prompt templates are provided here for reproducibility.
@@ 4.5 Multi-task Loss Balancing Strategy
During MCI model training, multi-task learning is performed jointly on change detection and change captioning, both using standard cross-entropy loss - applied to predicted change masks against ground truth annotations for detection, and to predicted sentences against ground truth descriptions for captioning. The two losses exhibit significant magnitude discrepancies, risking one task dominating the training signal. Additionally, conflicting gradients from the pixel-level detection and semantic captioning objectives can introduce destructive updates to shared representations during backpropagation.
To address this, we adopt a common multi-task loss balancing approach (Liu et al., 2024) by normalising losses to the same order of magnitude with gradients detached for stability, ensuring equal task contribution and reducing the risk of one task dominating the optimisation process:
@@ 4.6 Evaluation Metrics
To evaluate the performance of Forest-Chat’s supervised and zero-shot settings for binary change detection on RS imagery, the Mean Intersection over Union (mIoU) metric is employed, which quantifies segmentation quality by measuring the average overlap between predicted change masks and ground-truth masks across all classes:
@@
where $C$ is the number of classes (i.e. change (c), no change (nc)) for which metrics are calculated. For each class i, $TP_{i}$ are true positives, $FP_{i}$ false positives, and $FN_{i}$ false negatives for class $i$.
By analysing both the individual class performances through IoU, as well as mIoU, model predictions for the under-represented change class can be compared. Overall accuracy (OA) is naturally a poor performance indicator for this task due to the severe class imbalances present in all three datasets. Despite LEVIR-MCI-Trees containing three classes (i.e. road, building, background), performance metrics are calculated on the basis of change (c) and no change (nc), harmonising the results with those obtained for Forest-Change.
For evaluating the change captioning capabilities of FC-Supervised, FC-Zero-shot, and benchmarked models, common captioning evaluation metrics are applied for ease of comparison to the wider field. Additionally, BERTScore (Zhang et al., 2019) is employed to assess LLM-generated caption quality. The five metrics are summarised below:
@@
  • 5.
    BERTScore (Zhang et al., 2019): Computes semantic similarity between generated and reference captions using contextual embeddings from a pretrained transformer. By comparing cosine similarity between token embeddings rather than n-gram overlap, BERTScore captures paraphrasing and semantic equivalence, making it particularly suitable for evaluating LLM-generated captions.
@@ 5.2 Forest-Chat Supervised Performance
There remains a lack of available methods that simultaneously provide capabilities for joint change detection and captioning (CDC). Consequently, performance evaluations are done by comparing FC-Supervised to SOTA models for the two tasks separately on the LEVIR-MCI-Trees and Forest-Change datasets. Change3D (Zhu et al., 2025) is capable of both tasks, but is not implemented in a unified manner. BiFA (Zhang et al., 2024a) is selected for change detection, and Chg2Cap (Chang and Ghamisi, 2023) for change captioning. U-Net-based architectures remain popular for forest change segmentation (Nguyen et al., 2024), and segmentation tasks more generally (Corley et al., 2024). We therefore benchmark against a U-Net SiamDiff model employing a pretrained ResNet-50 backbone (He et al., 2016) with ImageNet (Deng et al., 2009) weights as a domain-specific baseline, following the implementation of Corley et al. (2024). Additionally, differences in difficulty between the two datasets are identified by comparing performance within a low-data forest change analysis scenario against an urban-focused one derived from a comprehensive source dataset.
Table 1 provides the performance comparisons for the benchmarked models on both aforementioned datasets. FC-Supervised is the strongest performer for change captioning overall, and is the best on LEVIR-MCI-Trees for change detection. On the Forest-Change dataset, FC-Supervised is marginally outperformed by BiFA Zhang et al. (2024a) for change class identification, with both models achieving similar overall performance. These results demonstrate strong generalisation capabilities across varying data domains for both change detection and captioning. The U-Net SiamDiff Corley et al. (2024) baseline was the worst performing change detection method on both datasets, particularly on the change class for Forest-Change. This highlights the value of designing architectures to explicitly process temporal information and task-aligned representations. The lightweight U-Net SiamDiff achieves the fastest change detection inference times, and the use of pretrained weights substantially improves performance compared to the non-pretrained Siamese Attention U-Net reported by Hewarathna et al. (2024) on the Forest-Change dataset (73.78% F1 vs 42.84% F1), from which our data is derived. Among the benchmarked methods, FC-Supervised is the slowest at captioning inference and the second slowest for change detection (see Table 2). This is expected because it performs joint multi-task reasoning with heavy cross-modal and cross-temporal interactions and multi-task attention. However, these inference times are not prohibitive for practical applications, especially given that only a single model is required to generate pixel-level and semantic-level information of changes within RS image pairs. Zero-shot inference times are discussed in Section 5.3.
@@
LEVIR-MCI-Trees U-Net SiamDiff 86.08 95.04 77.11 - - - - - - -
BiFA 87.54 95.63 79.45 - - - - - - -
@@
Forest-Change U-Net SiamDiff 64.42 95.71 33.14 - - - - - - -
BiFA 67.34 95.85 38.84 - - - - - - -
@@
Table 2: Inference time per iteration (s), averaged across LEVIR-MCI-Trees and Forest-Change datasets.
@@
Change Captioning FC-Zero-shot 3.34*
@@
Change Detection U-Net SiamDiff 0.12
Change3D 0.53
@@
FC-Zero-shotViT-B 16.07
FC-Zero-shotSAM2 23.28
FC-Zero-shotViT-H 31.88
*GPT-4o inference within FC-Zero-shot includes API communication overhead and is not directly comparable to local inference times.
@@
Overall change detection performance is considerably better for building and road detection than for deforestation, likely due to the pronounced class imbalances within Forest-Change, fuzzy edges at the borders of many deforestation patches, the uniqueness of deforestation patterns, and the relatively small dataset size. Within the Forest-Change dataset, a considerable percentage of change masks are comprised of numerous tiny patches that are challenging for models to detect. This resulted in all benchmarked CD models frequently predicting change masks that mirrored the deforestation patterns observed, but contained few overlaps with the ground-truth masks (see row 4 in Fig. 7). Models generally perform well on larger patches, but can still struggle with accurately predicting fuzzy boundaries. This corroborates the findings of Liu et al. (2024), who highlight that the segmentation performance of objects containing fewer than 400 pixels is low. Compared with deforestation detection, building change detection patterns and object geometries are more predictable, further raising the challenge posed by the Forest-Change dataset.
@@
Captioning performance evaluation is more nuanced. For the LEVIR-MCI-Trees dataset, the BLEU-1, BLEU-2, and CIDEr-D metrics are generally higher, while BLEU-3, BLEU-4, METEOR, and ROUGEL are better for Forest-Change. The observed results largely reflect the differences in annotation styles between the two datasets, as described in Section 3.2.3. Despite models frequently predicting deforestation severity accurately, they occasionally get confused about the location and characterisation of deforestation patches, and are unable to recognise geographic features within the imagery. To tackle this shortcoming, we investigate caption refinement through domain-aware zero-shot MLLMs in Section 5.4. Alternative domain-aware strategies, such as vocabulary expansion (Gao et al., 2024) or domain-adaptive fine-tuning (Guo et al., 2022), could also help address these limitations.
5.3 Forest-Chat Zero-shot Change Detection
In keeping with the previous section, FC-Zero-shot performance is evaluated on both the Forest-Change and LEVIR-MCI-Trees datasets. To explore the impact of different backbone architectures, three SAM-based configurations are evaluated: the default ViT-H backbone used in the Segment Any Change framework (Zheng et al., 2024), a smaller ViT-B backbone, and the more recent SAM2 (Ravi et al., 2024) architecture. These alternatives provide insight into the trade-offs between segmentation performance and inference efficiency for zero-shot change detection.
To optimise performance, a Bayesian hyperparameter search comprising 20 runs for each backbone was conducted to identify the most influential parameters. Following the Segment Any Change framework (Zheng et al., 2024), bi-temporal matching and feature normalisation are enabled throughout due to their consistently positive impact on change discrimination. The remaining hyperparameters are not tuned due to their minimal impact on mask generation. Points-per-side - which controls the spatial density of point prompts provided to SAM for mask proposal generation - is fixed at 16 as a practical trade-off between performance and inference time. Performing a full hyperparameter sweep on LEVIR-MCI-Trees is computationally prohibitive; therefore, the best configuration from Forest-Change is reused. These parameters mainly control proposal density and filtering heuristics, rather than dataset-specific semantics, so reusing them is unlikely to introduce systematic bias.